OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 9 −x − y = −8 −8 + y −8 − y 8 + y 8 − y

OpenStudy (anonymous):

Do you understand what the question asks for?

OpenStudy (anonymous):

If you do the answer is simple.

OpenStudy (anonymous):

This is the 2nd equation −x − y = −8 You want to use this to express x than use it to solve the 1st equation. For this you need to get an equation like x=something with only y in it.

OpenStudy (anonymous):

−x − y = −8 From this express it like x=.....

OpenStudy (anonymous):

Yes I understand what the question is asking for, what I don't understand is the outcome after O divide by -y.

OpenStudy (anonymous):

You dont have to divide by y

OpenStudy (anonymous):

(the equation divided by -y)

OpenStudy (anonymous):

Sorry -1

OpenStudy (anonymous):

there is no division! -x-y=-8 There is only subtraction in this equation

OpenStudy (anonymous):

take the -y first to the other side. So that the x bit is on one side and the y bit is on the other

OpenStudy (anonymous):

If you want to cancel a subtraction what do you do? What is the opposite of it?

OpenStudy (anonymous):

Addition

OpenStudy (anonymous):

-x-y+y=?

OpenStudy (anonymous):

-y=-8 + y

OpenStudy (anonymous):

-x=-8+y

OpenStudy (anonymous):

but you need +x not -x now what to do?

OpenStudy (anonymous):

What operation changes all the signs but leaves the numbers otherwise?

OpenStudy (anonymous):

Divide-x? Which would be x= 8-y?

OpenStudy (anonymous):

Nope. You multiply by -1

OpenStudy (anonymous):

If you multiply something with 1 you get the number itself. If you multiply by -1 that you get the numbers opposite

OpenStudy (anonymous):

And when you see x and y they are just numbers, unknown numbers

OpenStudy (anonymous):

So what happens with -x=-8+y if both sides are multiplied by -1?

OpenStudy (anonymous):

x=8-y?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Yea!!!!!

OpenStudy (anonymous):

Hope it cleared up a bit of confusion

OpenStudy (anonymous):

It really did thanks for taking out the time, I really appreciate it!